# What could go wrong? Discovering and describing failure modes in computer vision

### v1

@@ C.3 Pseudo-GT image-sentence relevance
To proceed with VQA, we first turn each sentence $s_{n}\in{\mathcal{S}}$ into a question $q_{n}$ such that the expected answer is yes or no. For example the sentence "An image taken at night" was turned into the prompt: “Was the image taken at night? Reply simply with ‘yes’ or ‘no’”. Then we provide VQA model with the image-question pairs $(x_{i},q_{n})$ and turn the yes/no answer into a binary score that is assigned to $\Gamma(x_{i},s_{n})$.
@@ Appendix D Design choices
In this section we detail our default parameter setting. We use Open-CLIP [17] trained on LAION-2B [39] as our default visual-textual embedding space, but one could use other VLM models, such as BLIP [24], etc. To generate the GT sentence set ${\mathcal{S}}^{*}_{\beta}$, we use $\beta=o*\omega_{\theta}^{\text{std}}$, where $\omega_{\theta}^{\text{std}}$ is the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$ with $o=0.2$ as default value. We set the number of hard and easy clusters to $C=15$ for large datasets (when we analyze the full set, namely urban scene segmentation and the unsupervised analyses on the NICO++{++}+ + datasets) and to $C=5$ for small sets, namely when we analyze the dataset for each class independently (per-class analyses of NICO++{++}+ + and ImageNet). We set the number of retained sentence per cluster for all methods to $K=3$.
@@
Figure 12: Five out of 15 hard clusters from NICO++{++}+ + 85 (unsupervised case) explained by different methods.
Appendix E Further qualitative analyses
First, we provide in Tab. 5 an extended version of Tab. 1 Then we show qualitative examples for NICO++{++}+ + 85 in Fig. 12, for ImageNet 1K in Figs. 13 and 14 and for WD2 in Fig. 15,
@@ Appendix E Further qualitative analyses
TopS SetDiff PDiff FPDiff
church       “taken at dusk       “taken in roof garden       “taken in a foggy weather
TopS SetDiff PDiff FPDiff
golden retriever       “a thumbnail image   “taken in hospital room   “a chow
TopS SetDiff PDiff FPDiff
typewriter keyboard       “with low contrast       “taken in clean room       “rule
computer keyboard       “rule       “with studio lighting       “hardware store
TopS SetDiff PDiff FPDiff
oxcart       “taken in medina       “oxcar       “oxcar
taken in a village       “taken in barndoor       “taken in slum       “taken in medina
taken in farm       “taken in hayfield       “taken in medina       “taken in a rice paddy
TopS SetDiff PDiff FPDiff
taken in bookstore       “taken in archive       “taken in slum       “taken in archive
bookshop       “taken in shed   “taken in porch   “taken in ticket booth
Figure 13: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false negative (not recognized) images from the correct class. From top to bottom we have the classes “church”, “golden retriever”, “computer keyboard”, “ox” and “bookshop”.
TopS SetDiff PDiff FPDiff
kit fox       “mongoose       “mongoose       “mongoose
grey fox       “taken in yard       “a grey fox       “a grey fox
red fox       “taken in wild field   “desert vegetation   “coyote
TopS SetDiff PDiff FPDiff
rock crab       “thumbnail image       “SD generated image       “taken in fishpond
Dungeness crab       “damaged image       “taken in galley       “Dungeness crab
fiddler crab       “taken in alcove   “taken at night   “taken in lagoon
TopS SetDiff PDiff FPDiff
stove       “taken in archive       “stove       “stove
taken in kitchen       “thumbnail image       “dishwasher       “dishwasher
TopS SetDiff PDiff FPDiff
zucchini       “taken in a farm       “vegetable garden       “vegetable garden
cucumber       “taken in a market outdoor       “taken in a farm       “taken in a farm
TopS SetDiff PDiff FPDiff
water snake       “green mamba       “thumbnail image       “thumbnail image
vine snake       “taken in arch   “taken in a cloudy weather   “green mamba
Figure 14: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false positives (images from other classes) images from the correct class. From top to bottom we have the classes “red fox”, “rock crab”, “microwave”, “cucumber” and “water snake”.
TopS SetDiff PDiff FPDiff
rainy weather      “rainy weather      “rainy weather      “rainy weather
vehicle on the road     “stormy weather      “stormy weather      “stormy weather
dull weather     “dust on the road     “dull weather     “dull weather
TopS     SetDiff     PDiff     FPDiff
rickshaw on the road      “sub-urban scene      “market      “rickshaw on the road
vehicle on the road     “dull weather      “restaurants      “taken in the morning
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “tunnel      “tunnel      “shadows on the road
car on the road     “guard-rail      “guard-rail      “light reflection on the road
motion blur     “dirt on the road     “snowy weather     “motion blur
TopS     SetDiff     PDiff      FPDiff
vehicle on the road      “snowy weather      “snowy weather      “snowy weather
clear weather     “rainy weather      “rainy weather      “rainy weather
snowy weather      “stormy weather     “stormy weather     “taken in winter
TopS SetDiff PDiff FPDiff
"vehicle on the road"      "mud on the road"      "mountain "      "mud on the road"
"obstacle on the road"      "taken in a windy weather"      "terrain"      "rocks on the road"
"jeep on the road"      "round-about scene"      "forest "      "jeep on the road"
TopS     SetDiff     PDiff     FPDiff
traffic      “taken at night      “taken at night      “taken at night
TopS SetDiff PDiff FPDiff
"tunnel"      "tunnel"      "tunnel"      "tunnel"
"vehicle on the road"      "taken at night"      "taken at night"      "light reflections on the road"
"car on the road"      "with high brightness"      "light reflections on the road"      "with motion blur"
TopS SetDiff PDiff FPDiff
"light reflections on the road"      "taken at night"      "taken at night"      "taken at night"
"taken at night"      "taken in the evening"      "taken at dusk"      "taken at dusk"
"vehicle on the road"      "taken at dusk"      "light reflections on the road"      "light reflections on the road"
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “taken in winter      “snowy weather      “taken in winter
clear weather     “snowy weather      “taken in winter      “dull weather
taken at daytime     “tunnel     “mountain     “snowy weather

### v2

@@ C.3 Pseudo-GT image-sentence relevance
To proceed with VQA, we first turn each sentence $s_{n}\in{\mathcal{S}}$ into a question $q_{n}$ such that the expected answer is yes or no. For example the sentence “An image taken at night” was turned into the prompt: “Was the image taken at night? Reply simply with ‘yes’ or ‘no’. ”. Then we provide VQA model with the image-question pairs $(x_{i},q_{n})$ and turn the yes/no answer into a binary score that is assigned to $\Gamma(x_{i},s_{n})$.
@@ Appendix D Design choices
In this section we recall and detail our default parameter setting. We use Open-CLIP [17] trained on LAION-2B [39] as our default visual-textual embedding space, but one could use other VLM models, such as BLIP [24], etc. To generate the GT sentence set ${\mathcal{S}}^{*}_{\beta}$, we use $\beta=o*\omega_{\theta}^{\text{std}}$, where $\omega_{\theta}^{\text{std}}$ is the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$ with $o=0.2$ as default value. We set the number of hard and easy clusters to $C=15$ for large datasets (when we analyze the full set, namely urban scene segmentation and the unsupervised analyses on the NICO++{++}+ + datasets) and to $C=5$ for small sets, namely when we analyze the dataset for each class independently (per-class analyses of NICO++{++}+ + and ImageNet). We set the number of retained sentences per cluster for all methods to $K=3$.
TopS SetDiff PDiff FPDiff
church       “taken at dusk       “taken in roof garden       “taken in a foggy weather
TopS SetDiff PDiff FPDiff
golden retriever       “a thumbnail image   “taken in hospital room   “a chow
TopS SetDiff PDiff FPDiff
typewriter keyboard       “with low contrast       “taken in clean room       “rule
computer keyboard       “rule       “with studio lighting       “hardware store
TopS SetDiff PDiff FPDiff
oxcart       “taken in medina       “oxcar       “oxcar
taken in a village       “taken in barndoor       “taken in slum       “taken in medina
taken in farm       “taken in hayfield       “taken in medina       “taken in a rice paddy
TopS SetDiff PDiff FPDiff
taken in bookstore       “taken in archive       “taken in slum       “taken in archive
bookshop       “taken in shed   “taken in porch   “taken in ticket booth
Figure 12: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false negative (images from the current class not recognized). From top to bottom we have the classes “church”, “golden retriever”, “computer keyboard”, “ox” and “bookshop”.
@@
Appendix E Further qualitative analyses
[…]
TopS SetDiff PDiff FPDiff
kit fox       “mongoose       “mongoose       “mongoose
grey fox       “taken in yard       “a grey fox       “a grey fox
red fox       “taken in wild field   “desert vegetation   “coyote
TopS SetDiff PDiff FPDiff
rock crab       “thumbnail image       “SD generated image       “taken in fishpond
Dungeness crab       “damaged image       “taken in galley       “Dungeness crab
fiddler crab       “taken in alcove   “taken at night   “taken in lagoon
TopS SetDiff PDiff FPDiff
stove       “taken in archive       “stove       “stove
taken in kitchen       “thumbnail image       “dishwasher       “dishwasher
TopS SetDiff PDiff FPDiff
zucchini       “taken in a farm       “vegetable garden       “vegetable garden
cucumber       “taken in a market outdoor       “taken in a farm       “taken in a farm
TopS SetDiff PDiff FPDiff
water snake       “green mamba       “thumbnail image       “thumbnail image
vine snake       “taken in arch   “taken in a cloudy weather   “green mamba
Figure 13: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false positive (images from other classes wrongly labeled with the current class name). From top to bottom we have the classes “red fox”, “rock crab”, “microwave”, “cucumber” and “water snake”.
@@ Appendix E Further qualitative analyses
TopS   SetDiff   PDiff   FPDiff
motorcycle in the rocks       “motorcycle in the rocks       “motorcycle in the rocks       “motorcycle in the rocks
motorcycle   “motorcycle       “bicycle in the rocks       “bicycle in the rocks
landways   “landways   “car in the rocks   “car in the rocks
TopS   SetDiff   PDiff   FPDiff
helicopter in the rocks       “hot air balloon in the rocks       “hot air balloon in the rocks       “hot air balloon in the rocks
helicopter   “airways       “bicycle in the rocks       “helicopter in the rocks
helicopter taken at sunset   “airplane in the water   “ train in the rocks   “airplane in the rocks
[…]
TopS SetDiff PDiff FPDiff
rainy weather      “rainy weather      “rainy weather      “rainy weather
vehicle on the road     “stormy weather      “stormy weather      “stormy weather
dull weather     “dust on the road     “dull weather     “dull weather
TopS     SetDiff     PDiff     FPDiff
rickshaw on the road      “sub-urban scene      “market      “rickshaw on the road
vehicle on the road     “dull weather      “restaurants      “taken in the morning
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “tunnel      “tunnel      “shadows on the road
car on the road     “guard-rail      “guard-rail      “light reflection on the road
motion blur     “dirt on the road     “snowy weather     “motion blur
TopS     SetDiff     PDiff      FPDiff
vehicle on the road      “snowy weather      “snowy weather      “snowy weather
clear weather     “rainy weather      “rainy weather      “rainy weather
snowy weather      “stormy weather     “stormy weather     “taken in winter
TopS SetDiff PDiff FPDiff
vehicle on the road      “mud on the road      “mountain      “mud on the road
obstacle on the road      “taken in a windy weather      “terrain      “rocks on the road
jeep on the road      “round-about scene      “forest      “jeep on the road
TopS     SetDiff     PDiff     FPDiff
traffic      “taken at night      “taken at night      “taken at night
TopS SetDiff PDiff FPDiff
tunnel      “tunnel      “tunnel      “tunnel
vehicle on the road      “taken at night      “taken at night      “light reflections on the road
car on the road      “with high brightness      “light reflections on the road      “with motion blur
TopS SetDiff PDiff FPDiff
light reflections on the road      “taken at night      “taken at night      “taken at night
taken at night      “taken in the evening      “taken at dusk      “taken at dusk
vehicle on the road      “taken at dusk      “light reflections on the road      “light reflections on the road
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “taken in winter      “snowy weather      “taken in winter
clear weather     “snowy weather      “taken in winter      “dull weather
taken at daytime     “tunnel     “mountain     “snowy weather